OpenStudy (anonymous):

Infinite sum of the sequence (2^(1/n)-1) Prove Divergence or Convergence

OpenStudy (anonymous):

check whether the common ratio is |r| < 1 if < 1, it's convergent. Otherwise, it's divergent

OpenStudy (anonymous):

See, that's the thing, you can't do that with 2^(1/n), ratio test doesn't solve anything.

OpenStudy (zarkon):

limit comparison with 1/n

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} n2^{1/n} - n\] isn't that an indeterminate form

OpenStudy (zarkon):

\[\lim_{n\to\infty}\frac{2^{1/n}-1}{1/n}\] use L'Hospitals rule

OpenStudy (anonymous):

\[-\frac{ 2^{1/n} \log 2 }{ n ^{2} } / n ^{-2}\] like so?

OpenStudy (zarkon):

\[\lim_{n\to\infty}\frac{2^{1/n}-1}{1/n}=\lim_{n\to\infty}\frac{2^{1/n}\left(\frac{-1}{n^2}\right)\ln(2)}{\frac{-1}{n^2}}=\lim_{n\to\infty}2^{1/n}\ln(2)\]

OpenStudy (anonymous):

That goes to ln(2), which is less than 1, yes? Which is a constant... meaning that it converges because of definition of LCT?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

ln(2) is a positive finite number so \[\sum_{n=1}^\infty(2^{1/n}-1)\] does what \[\sum_{n=1}^\infty\frac{1}{n}\] does

OpenStudy (anonymous):

Yes, I understand. So lim of An/Bn goes to a constant, and An does what Bn does. Bn diverges, so An must diverge, yes?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (zarkon):

np